OpenStudy (anonymous):

Stan is making a trail mix that has a combination of peanuts and raisins. The mix is 3 parts peanuts and 1 part raisins. How many cups of raisins does he need to use for a 16-cup recipe? 5cups 4 cups cups 64 cups

OpenStudy (anonymous):

mixture is=peanuts:raisins=3:1 so in (3+1)cup i.e 4 cup there will be 1 cup raisin so for 16 cup recipe raisin required is=(1/4)*16=16/4=4 cup.

OpenStudy (anonymous):

i do not get it

OpenStudy (anonymous):

its a ratio 3 part to 1 part

OpenStudy (anonymous):

3:1

OpenStudy (anonymous):

3 peanuts 1 raisins

OpenStudy (anonymous):

you cant just give me the answer

OpenStudy (anonymous):

yeh i didnt

OpenStudy (anonymous):

the answer is 4 cups

OpenStudy (anonymous):

u would do 16 / 4

OpenStudy (phi):

the thing is get this idea: 3 parts peanuts and 1 part raisins. imagine each "part" is 1 cup, so you have 3 cups of peanuts and 1 cup of raisins. First question: how many cups total of stuff do you have ?

OpenStudy (anonymous):

it would be 4

OpenStudy (anonymous):

Which of the following items has the highest unit price? 3 for $2.19 4 for $3.00 5 for $3.50 $0.72 each

OpenStudy (anonymous):

then u would do 16 / 4

OpenStudy (anonymous):

plz post as seperate problem

OpenStudy (anonymous):

jacob you are the best on the web site

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ok 1st do .72 x 3 if it goes over 2.19 then it would be highest as of those 2

OpenStudy (anonymous):

then do 4 and then 5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would the answer be b.

OpenStudy (anonymous):

i dont have anything with me right now 1 sec

OpenStudy (anonymous):

yes try that

OpenStudy (anonymous):

A bookstore marks up the cost of a book from $6 to $10. What was the percent increase? 66.7% 34% 16% 2%

OpenStudy (anonymous):

the price increased by $4 therefore = 4/6*100 = 66.67 % increase

OpenStudy (anonymous):

a

OpenStudy (anonymous):

the answer is a.

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

ΔMNO ~ ΔPQR What is the length of x? 2 2 1 1

OpenStudy (anonymous):

i cant answer this one

OpenStudy (anonymous):

it has questions covering the question

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

On a map, 2 centimeters represents 100 kilometers. How much of a centimeter represents 20 kilometers? cm cm cm cm

OpenStudy (anonymous):

1/2 1/10 1/5 2/5

OpenStudy (anonymous):

by

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

srry if i dont reply its cause im doing stuff

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

@hunter.glover

OpenStudy (anonymous):

u get dis ?